OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates. ordered pair negative 5 comma 5 pi divided by 3

OpenStudy (anonymous):

the ordered pair is like \[(-5, 5\pi/3)\]

OpenStudy (misty1212):

no no that is backwards!!

OpenStudy (misty1212):

\[x=r\cos(\theta), y = r\sin(\theta)\]

OpenStudy (anonymous):

so x=-2.5

OpenStudy (misty1212):



OpenStudy (misty1212):

because of the \(-5\) you have to go in the other direction

OpenStudy (misty1212):

yeah i think \(-2.5\) is right

OpenStudy (misty1212):

same as \[5\cos(\frac{2\pi}{3})=5\times (-\frac{1}{2})=-2.5\]

OpenStudy (misty1212):

the y values should be \[\frac{5\sqrt3}{2}\] if i am not mistaken

OpenStudy (anonymous):

okay cool! thats one of my answers! thank you!

OpenStudy (misty1212):

lol i sure hope it is right!

OpenStudy (anonymous):

my last question would be this:

OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (3, 3) with 0° ≤ θ < 360°

OpenStudy (misty1212):

hmm i got the question marks let me refresh

OpenStudy (misty1212):

oh ok

OpenStudy (misty1212):

now we have switched to degrees i see

OpenStudy (misty1212):

pretty clear that the angle is \(45^\circ\) right?

OpenStudy (anonymous):

\[0 degrees \le \theta \le 360degrees\]

OpenStudy (anonymous):

im not sure, this one has me really confused

OpenStudy (misty1212):



OpenStudy (misty1212):

right 3, up 3 angle is \(45^\circ\)

OpenStudy (misty1212):

the length is \[\sqrt{3^2+3^2}=\sqrt{2\times 3^2}=3\sqrt2\]

OpenStudy (misty1212):

so one answer could be \[(3\sqrt2,45^\circ)\]

OpenStudy (anonymous):

okay and then make the first one negative right?

OpenStudy (misty1212):

yes if you make the angle \(125^\circ\)

OpenStudy (misty1212):

is that a choice?

OpenStudy (anonymous):

so \[(-3\sqrt{2}, 225)\]

OpenStudy (misty1212):

\[(-3\sqrt2,225^\circ)\]

OpenStudy (misty1212):

bingo!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (anonymous):

and i lied, i actually have one last question that i thought i understood, but i managed to confuse myself some more!